How would you find x and y by elimination of substituting this problem?
-3x+2y=-6
-2x+y=6

Answer by Ashley1516(4)

OK.....To solve this problem by substitution, you first need to get either x or y by itself on one side of the equal sign. The y in eq.2(-2x+y=6) is the easiest to work with. Change the -2x to positive 2x and add to itself and 6. Like so:
-2x+y=6
+2x +2x
-----------
y=2x+6 Now you know what y is and u can substitute it into eq.1(-3x+2y=-6)! Like so:
-3x+2(2x+6)=-6 Now solve for x:
-3x+4x+12=-6
x+12=-6
-12 -12
----------
*x=-18* Now u have x!! Plug x into eq.2 to find y:
-2(-18)+y=6
36+y=6
-36 -36
---------
*y=-30* You now have y!!
YOUR ANSWER IS (x,y):
*-18,-30*
Always make sure to check your answer!
Just plug in your y and x into both equations:
-3x+2y=-6(eq.1)
-3(-18)+2(-30)=-6
54+(-60)=-6
-6=-6 CORRECT!
-2x+y=6
-2(-18)+(-30)=6
36+(-30)=6
6=6 CORRECT!
Well, I really hope this helps and you can understand it!
~aShLeY~